OpenStudy (anonymous):

A racquetball club charges $10 for a one-month trial membership. After the trial month, the regular membership fee is $12 per month. Write an equation that gives the total cost of a membership C as a function of the length of membership m (in months). (NOTE: Your equation must use variables m and C and be in slope-intercept form, C=...)

OpenStudy (anonymous):

y=12(m-1) +10 and subtract 10 over (y-10)=12(m-1)

OpenStudy (anonymous):

thank you very much sir!

OpenStudy (anonymous):

wait. It has to be in y=mx+b form